bella2025:

a theater sells tickets for a concert. Tickets for lower level seats sell for $35 each and tickets for upper level seats sell for $25 each. The theater sells 350 tickets for $10.250 How many tickets of each type were sold

bella2025:

yeah I dont know what im doing

dukedadon:

Are there any answer choices?

Sailor:

Ok. We can set up equations for the two different types of tickets. Tickets for the lower level seats: \[35x\] Tickets for the upper level seats:\[25y\] These are going to eventually be in the same equation, but the x and the y shows that they are two different things.

Sailor:

Is 10.250 10 dollars and 25 cents or is is something else?

bella2025:

@dukedadon wrote:
Are there any answer choices?
no

bella2025:

@sailor wrote:
Is 10.250 10 dollars and 25 cents or is is something else?
no thats 10 dollars and 25 cents

Sailor:

@bella2025 wrote:
@sailor wrote:
Is 10.250 10 dollars and 25 cents or is is something else?
no thats 10 dollars and 25 cents
Ok.

Sailor:

Are you sure it's not 10,250?

Sailor:

Because if we set up an equation like 35x+25y=10.250, it just wouldn't work.

Sock:

Question, what grade are you in?

bella2025:

@sock wrote:
Question, what grade are you in?
9th

bella2025:

@sailor wrote:
Because if we set up an equation like 35x+25y=10.250, it just wouldn't work.
ok so then Ig its 10,250

Sock:

@bella2025 wrote:
@sailor wrote:
Because if we set up an equation like 35x+25y=10.250, it just wouldn't work.
ok so then Ig its 10,250
can u take a ss of the question so we can see?

bella2025:

@sock wrote:
@bella2025 wrote:
@sailor wrote:
Because if we set up an equation like 35x+25y=10.250, it just wouldn't work.
ok so then Ig its 10,250
can u take a ss of the question so we can see?
Its on paper and my camera dont work

Sailor:

Ok. Well then the first equation would be $35x + $25y = $10,250 and the second equation would be x + y = 350.

bella2025:

@sailor wrote:
Ok. Well then the first equation would be $35x + $25y = $10,250 and the second equation would be x + y = 350.
ok so then what would I do from there

Sailor:

You would use the substitution method.

bella2025:

@sailor wrote:
You would use the substitution method.
plug 1 in for y and x

Sailor:

Ok. On the second equation, you would get y on the other side. Then, you would plug 350-x into the equation $35x + $25y = $10,250. \[35x + 25 (350 - x) = $10,250\] \[35x + $8750 - 25x = $10,250\] \[10x + $8750 = $10,250\] Subtract 8,750 from both sides \[10x = 1,500\] Divide by 10 on each side. You should get x = 150.

Sailor:

You would then substitute that into the equation of y = 350 - x. \[y=350-x\] \[y = 350 - 150\] \[y=200\]

Sailor:

Therefore, x=150 and y=200. So, there were 150 tickets for the lower seats, and 200 tickets for the upper level seats.

Sailor:

@bella2025 Hope this helps you out.

bella2025:

@sailor wrote:
@bella2025 Hope this helps you out.
thank you ive just had so much going on so for me not to answer im sorry

Sailor:

@bella2025 wrote:
@sailor wrote:
@bella2025 Hope this helps you out.
thank you ive just had so much going on so for me not to answer im sorry
You're welcome.